OpenStudy (anonymous):

Please help. There are 29 people at a Mardi Gras party. 11speak French, 24 speak English, and 3 speak neither French nor English. How many people at the party speak both English and French?

OpenStudy (swissgirl):

There are a total of 29 people at the party Now there are 11speak french, 24 speak English and 3 that speak neither 11+24+3=38 There are only 29 people at the party so some must speak french and English 38-29=9 There are 9 people that speak both languages

OpenStudy (anonymous):

but would you count out the three people who speak neither?

OpenStudy (swissgirl):

Nope since they are part of the 29 people at the party What you can do is 29-3=26 There are 26 people that speak french or english or both There are 24 that speak English and 11 that speak french 24+11=35 35-26=9

OpenStudy (anonymous):

Okay. Thank You!